# Solving the Hubbard model with Neural Quantum States

### v1

@@ Abstract
The rapid development of neural quantum states (NQS) has established it as a promising framework for studying quantum many-body systems. In this work, by leveraging the cutting-edge transformer-based architectures and developing highly efficient optimization algorithms, we achieve the state-of-the-art results for the doped two-dimensional (2D) Hubbard model, arguably the minimum model for high-Tc superconductivity. Interestingly, we find different attention heads in the NQS ansatz can directly encode correlations at different scales, making it capable of capturing long-range correlations and entanglements in strongly correlated systems. With these advances, we establish the half-filled stripe in the ground state of 2D Hubbard model with the next nearest neighboring hoppings, consistent with experimental observations in cuprates. Our work establishes NQS as a powerful tool for solving challenging many-fermions systems.
@@
The Hubbard model   (?, ?, ?) is the iconic model for studying correlation effect in many-electron systems. Though its Hamiltonian is simple, the physics it can harbor is extremely rich, including quantum magnetism, Mott insulator, charge and spin density waves and so on. Following the discovery of cuprate high-Tc superconductors, the two-dimensional (2D) Hubbard model has attracted increasing attention due to its potential connection to the microscopic mechanism of high-Tc superconductivity  (?, ?). The study of the 2D Hubbard model relies mainly on numerical many-body methods due to the lack of analytic solution. In the past few decades, a variety of many-body methods were developed, including density matrix renormalization group (DMRG)  (?) and its higher dimension generalization in the langurage of tensor network states  (?, ?), such as projected entangled pair states (PEPS)  (?), quantum embedding methods such as dynamical mean-field theory (DMFT) (?), density matrix embedding theory (DMET) (?), and different flavors of quantum Monte Carlo approaches  (?, ?, ?) and so on. In the past, different methods usually give inconsistent results due to the limitation in each method  (?). In recent years, large scale cross-validation has been performed among different methods and consensus has began to reach about the ground state of the doped 2D Hubbard model, with  (?) and without the next nearest neighboring hoppings  (?).
@@
Figure 1: Application, architecture, optimization, and performance of Neural Quantum States (NQS) on the Hubbard model. (A) Guided by the variational principle, NQS are powerful and versatile, applicable to a wide range of quantum systems, including ab-initio quantum chemistry, spin models, and correlated fermionic models. (B) Various neural network architectures can be used to construct NQS, such as Restricted Boltzmann Machines (RBM) (?), Multi-Layer Perceptrons (MLP)  (?, ?, ?, ?), and Transformers. These networks are optimized using methods inspired by quantum information, including Stochastic Reconfiguration (SR)  (?, ?, ?), Min-SR  (?), SPRING  (?), and the advanced MARCH method proposed in this work. (C) A performance comparison of different NQS methods on the 2D Hubbard model, plotting simulation accuracy against system size. We benchmark the accuracy on systems where exact solutions are available (see Supplementary Section 5.5 for details). This work achieves higher accuracy on larger system sizes than previously reported in other notable NQS studies, demonstrating a significant advancement in the field.
@@
On another front, researchers are also striving to develop more accurate and efficient new methods, aiming to provide more reliable results for strongly correlated systems like Hubbard model to unveil the underlying exotic physics. In recent years, benefiting from the rapid development in the fields of neural networks and machine learning, Neural Quantum States (NQS) have emerged as a promising new approach for solving quantum many-body systems. Significant progress has been made in the development of advanced NQS methods in quantum systems since their inception by Carleo et al.  (?), which have markedly improved the accuracy of NQS approaches, establishing them as a powerful framework for investigating systems in quantum chemistry  (?, ?, ?) and interacting spin models  (?, ?), as shown in Figure 1(A). In light of these significant achievements, a crucial question arises: how effective is this new paradigm for the study of the more challenging strongly correlated fermionic lattice models?
@@
In this direction, extensive effort has been made to investigate the utilization of NQS methods in the context of the Hubbard model, particularly in the most challenging 2D lightly doped cases. Early attempt of applying NQS approaches to the 2D Hubbard model (?) faced challenges in accurately representing fermionic nodal structures. Subsequent work by Luo et al. (?) introduced a neural network backflow ansatz that substantially improved the representation power of the wavefunction, demonstrating its effectiveness on systems as large as $12\times 8$. More recently in 2022, Robledo et al.  (?) achieved highly accurate results for various geometries of the Hubbard model up to $8\times 8$ by combining Hidden-Fermion Determinant States (HFDS) with Stochastic Reconfiguration (SR)  (?, ?, ?). Despite these notable progress, scaling NQS simulations to even larger, more physically relevant parameter regimes is hindered by network architectural limitations and optimization barriers (?, ?).
@@
In this work, we make significant progress in expanding the scalability and accuracy of the NQS approach for solving the Hubbard model with two key innovations as shown in Figure 1(B). First, a novel NQS architecture is proposed, which leverages the self-attention mechanism  (?), a technique proven to be very effective in scaling up  (?). The built-in long-range structure make the NQS ansatz different from tensor network states which consists of only local tensors. This ensures that the NQS can precisely capture long-range correlations and entanglements in strongly correlated systems and can accurately handle larger systems, which is infeasible for other neural network architectures like Restricted Boltzmann Machine (RBM)  (?) and Multi-Layer Perceptron (MLP)  (?, ?, ?, ?). Second, we propose an enhanced optimization algorithm, termed Moment-Adaptive ReConfiguration Heuristic (MARCH), designed to achieve substantially faster and more stable convergence than previous optimization algorithms such as SR  (?, ?, ?), Min-SR  (?) and SPRING  (?), especially for large and frustrated systems.
@@
With these architectural and algorithmic advancements, we achieve state-of-the-art variational ground state energy of the 2D Hubbard model at unprecedented sizes up to $16\times 16$, for both open and periodic boundary conditions, as shown in Figure 1(C). As a result, we confirm filled stripe ground state in the pure Hubbard model and identify a half-filled stripe ground state with the next nearest neighboring hopping $t^{\prime}=-0.2$, consistent with experimental observations in cuprate   (?). Moreover, in the $t^{\prime}=-0.2$ case, we find the stripe is horizontal, i.e., stripe arrangements prefer the longer side of the geometry. This is quite different from previous calculations on cylinders  (?), indicating the effectiveness of our method in eliminating the boundary effects. We also have a careful analysis of the structure of our NQS ansatz and find that different attention heads can directly encode correlations at different scales after optimization, making it capable to capture long-range correlations and entanglements in the Hubbard model. Our developments in architectures and optimization strategies establish NQS as a powerful and transformative paradigm for solving the Hubbard model. This framework extends readily to a broader range of strongly correlated systems, offering a new lens to explore new exotic quantum many-body states resulting from strong correlations.
@@
Here $\hat{c}_{i\sigma}^{\dagger}$ $(\hat{c}_{j\sigma})$ is the creation (annihilation) operator for an electron with spin $\sigma$ at site $i(j)$, $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$ is the particle number operator. The terms $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote pairs of nearest neighboring and next nearest neighboring sites, respectively. We study the 2D Hubbard model defined on square lattice with size $L_{x}\times L_{y}$ (with $L_{x}\geq L_{y}$). Both open boundary conditions (OBC) and periodic boundary conditions (PBC) are adopted in this work. We set the nearest neighboring hopping amplitude $t=1$ as the unit of energy. The on-site Coulomb repulsion is fixed at $U=8$, a value relevant to cuprates  (?, ?). We consider both the pure Hubbard case ($t^{\prime}=0$) and case with the next nearest neighboring hoppings ($t^{\prime}=-0.2$). In this work, we focus on the lightly doped regime with hole concentration of $\delta=1/8$, which is the most intriguing region in both the 2D Hubbard model and in cuprates.
@@
Figure 2 presents an overview of our NQS algorithm. As shown in Figure 2(A), we employ a transformer-based NQS, parameterized by $\theta$, to represent the amplitude $\psi_{\theta}(\mathbf{n})$ associated with each configuration |$\ket{\mathbf{n}}$. Our transformer-based NQS assigns an unique high-dimensional embedding vector to each site $i$. These embeddings are subsequently processed through a series of multi-head attention blocks, the core component of modern large language models  (?). The attention block is defined as:
@@
where $Q,K,V\in\mathbb{R}^{N\times d_{H}}$ are feature matrices produced by neural networks. Here, $N$ represents the number of sites in the Hubbard model, $d_{H}$ is the hidden dimension of the attention block. The normalized attention scores $S_{\text{attn}}$, given by $\text{softmax}(QK^{\top}/\sqrt{d_{H}})$, allow the block to aggregate information from all sites by assigning weights. The $i$-th row of these matrices corresponds to the $i$-th site in the Hubbard model, arranged in dictionary order. Thus, the attention block can aggregate information across all sites, capturing complex, long-range interactions and correlations, which is crucial for modeling strongly correlated systems like the Hubbard model. The multi-head attention block employs distinct feature matrices to capture diverse correlation patterns in the system. After self-attention layers, a linear layer is employed to produce backflow orbitals $\mathbf{M}(\mathbf{n})$. The amplitude is then derived as the determinant of orbitals occupied by electrons.
In our work, we aim to find the ground state by minimizing the total energy of the wavefunction. In previous NQS studies (?), SR-based optimizers have been used to minimize the energy. These optimizers effectively simulate the imaginary time evolution, enhancing the convergence rate over standard first-order optimizer by exploiting the quantum geometric tensor  (?). One particular method is the SPRING algorithm  (?), which leverages the history of gradients to accelerate convergence and smooth out the optimization path. Building on this, we introduce the MARCH algorithm. MARCH enhances SPRING by also incorporating an estimate of the second moment of the gradients, drawing a direct parallel to how the popular Adam optimizer  (?) improves upon Momentum  (?). This addition of second-moment information is crucial for navigating the complex and high-dimensional energy landscapes typical existing in quantum systems. It provides an estimate of the gradient’s variance, allowing MARCH to adapt the learning rate for each parameter individually. This adaptive capability offers a significant advantage, particularly when encountering challenging regions like saddle points or flat plateaus where the gradient is small. While a momentum-based method like SPRING might overshoot, MARCH can use the second-moment information to maintain progress and “march” confidently out of these areas. As illustrated in Figure 2(B), this results in a more direct and efficient path to the energy minimum, leading to faster and more robust convergence.
@@
We then benchmark the new NQS method on doped pure Hubbard model under OBC with varying sizes $L_{x}$ and $L_{y}$. Direct comparisons of ground-state energies with state-of-the-art DMRG and PEPS  (?) results are presented in Figure 2(C) (with Hartree-Fock energies as reference). In narrow systems such as $L_{y}\leq 6$, the DMRG energy is the lowest. However, when reaching the 2D limit by increasing the width of the system, where DMRG fails to capture the area law of entanglement entropy, the NQS results surpass DMRG. Noticeably, in the entire range of $L_{y}$ we studied, NQS achieves lower variational energies than those reported by state-of-the-art PEPS with bond dimension $D\geq 20$  (?).
@@
Figure 3: Visualization of the attention mechanism, showing how a lattice site attends to another target site in the 2D pure Hubbard model for lattice sizes of $8\times 8$, $10\times 10$, and $12\times 12$ at half filling. The four attention heads, which are equivalent and have been rearranged by pattern for clarity, each corresponds to a specific physical aspect: short-range correlations, nearest-neighbor hopping, antiferromagnetic correlations, and complex and delocalized correlation in the ground state (details and similar results for the doped cases can be found in Figure S26 to Figure S35 in the Supplementary Materials).
@@
To gain deeper insight into how the transformer architecture captures the intricate physics of the Hubbard model, we examine the attention patterns learned by the network. In our transformer ansatz, the heads after optimization are found to specialize to capture the correlations in different scales present in the ground-state wave function. This emergent specialization is analogous to observations in fields like natural language processing, where different attention heads in a transformer learn distinct, interpretable patterns such as syntactical dependencies  (?). In Figure 3, we visualize the normalized attention scores $S_{\text{attn}}$ for the half-filling systems, which quantify the learned pairwise importance between sites, from the first layer for representative systems. Similar results for the doped cases can be found in Figures S26-S35 in the Supplementary Materials. The results are striking: the first head clearly learns to prioritize local interactions, with $S_{\text{attn}}$ that decay with the distance between sites, capturing short-range correlations. The second head specializes in nearest-neighbor hopping, with its attention focusing almost exclusively on the four adjacent sites, which is a fundamental process in the Hubbard model. The third head successfully identifies the emergent antiferromagnetic order, assigning high attention scores in a checkerboard pattern characteristic of this magnetic correlation. The final head appears to capture more complex, delocalized phase information, exhibiting a sophisticated pattern that is crucial for accurately representing the full complexity of the wave function. This specialization demonstrates the remarkable ability of the transformer’s attention mechanism to autonomously discover and disentangle the fundamental physical correlations governing the system.
@@
Figure 4: Hole and spin density distributions in the ground state for different lattice sizes with $U=8$ and hole doping $\delta=1/8$ under PBC in the Hubbard model. The magnitudes of the spin density are represented by the sizes of the arrows while the direction is denoted by the direction of the arrows. Hole density is depicted using a color scale. Stripes can be clearly seen in all three systems. (A) Results for $16\times 16$ pure Hubbard model where stripe with $\lambda=8$ is observed. (B) Results for $16\times 12$ system with $t^{\prime}=-0.2$ where stripe with $\lambda=4$ is observed. (C) Results for $32\times 8$ system with $t^{\prime}=-0.2$ where stripe with $\lambda=4$ is observed. Notice that the stripes in the ground state for $t^{\prime}=-0.2$ in (B) and (C) are horizontal even though both vertical and horizontal stripes are allowed in the systems. More results can be found in the supplementary materials.
The ground state of the lightly doped pure ($t^{\prime}=0$) 2D Hubbard model was established as the filled stripe before  (?). Recently in  (?), a crossover of the stripe direction was observed with the increase of system widths for systems under OBC, indicating the existence of large finite size effect in systems with open boundaries. We performed systematic calculations of the pure Hubbard model for different sizes under both OBC and PBC. For systems with OBC, we find stripe states that match well with the results obtained from PEPS but with lower energies (see Figure 2(C)). In particular, our results also demonstrate a crossover between horizontal and vertical stripes as a function of system width (from $L_{y}=8$ to $L_{y}=12$, see Figures S7-S8 in the Supplementary Materials). But when we switch to PBC, the crossover vanishes (see Figures S11-S12 in the Supplementary Materials), indicating the crossover is a finite size and open boundary effect. The ability to handle systems under PBC with NQS can mitigate the finite size effect, allowing to obtain more reliable results in the thermodynamic limit. By studying systems under PBC, we observe a robust, filled stripe phase with wavelength of $\lambda=8$ as shown in Figure 4(A), a result that is consistent with previous findings  (?). While simulations on smaller $8\times 8$ and $12\times 12$ lattices (see Figures S10S13 in the Supplementary Materials) shows clear finite-size effects due to their incommensurability with the wavelength $\lambda=8$.
@@
It was found that the doped pure Hubbard model doesn’t host superconductivity  (?). To pursuit superconducting ground state, the simplest modification is to include the next nearest neighboring hopping ($t^{\prime}$) term. On one hand, a finite $t^{\prime}$ can account for the electron and hole doping asymmetry in the phase diagram of cuprates. On the other hand, electronic structure calculation  (?, ?) show a finite $t^{\prime}\approx-0.2$ term indeed exists in the one band effective Hamiltonian of cuprates. A recent study  (?) found that the phase diagram of Hubbard model with $t^{\prime}=-0.2$ agrees well with the cuprates. In  (?), fractional filling stripe is found to coexist with the d-wave superconductivity in the hole doping region. The $t^{\prime}$ term frustrates the anti-ferromagnetic order in the parent compound and make the doped stripe state fragile. Importantly, large finite size and boundary effect was found for the wave-length of the stripe state in the hole doping region, which fluctuates in the vicinity of $\lambda=4$  (?).
We perform systematic calculations for systems under PBC with $t^{\prime}=-0.2$. With PBC, we can get rid of the boundary effect. Our calculations on large sizes show the ground state for the $t^{\prime}=-0.2$ Hubbard model is clearly half-filled stripe, i.e., with wavelength $\lambda=4$, as shown in Figure 4 (B). These results indicate that the previous found fluctuating wavelengths for stripe may results from the finite size and boundary conditions. The $\lambda=4$ stripe agree well with the experimental observation of cuprates  (?), providing further evidence to support the relevance of the $t^{\prime}$ Hubbard model to cuprate.

### v2

@@ Abstract
The rapid development of neural quantum states (NQS) has established it as a promising framework for studying quantum many-body systems. In this work, by leveraging the cutting-edge transformer-based architectures and developing highly efficient optimization algorithms, we achieve the state-of-the-art results for the doped two-dimensional (2D) Hubbard model, arguably the minimum model for high-Tc superconductivity. Interestingly, we find different attention heads in the NQS ansatz can directly encode correlations at different scales, making it capable of capturing long-range correlations and entanglements in strongly correlated systems. With these advances, we establish the half-filled stripe in the ground state of 2D Hubbard model with the next nearest neighboring hopping, consistent with experimental observations in cuprates. Our work establishes NQS as a powerful tool for solving challenging many-fermions systems.
@@
The Hubbard model  (?, ?, ?) is the iconic model for studying the correlation effect in many-electron systems. Though its Hamiltonian is simple, the physics it can harbor is extremely rich, including quantum magnetism, Mott insulator, charge and spin density waves, and so on. Following the discovery of cuprate high-Tc superconductors, the two-dimensional (2D) Hubbard model has attracted increasing attention due to its potential connection to the microscopic mechanism of high-Tc superconductivity  (?, ?). The study of the 2D Hubbard model relies mainly on numerical many-body methods due to the lack of analytic solutions. In the past few decades, a variety of many-body methods were developed, including density matrix renormalization group (DMRG)  (?) and its higher dimension generalization in the language of tensor network states  (?, ?), such as projected entangled pair states (PEPS)  (?), quantum embedding methods such as dynamical mean-field theory (DMFT) (?), density matrix embedding theory (DMET) (?), and different flavors of quantum Monte Carlo approaches  (?, ?, ?) and so on. In the past, different methods usually gave inconsistent results due to the limitations of each method  (?). In recent years, large scale cross-validation has been performed among different methods, and consensus has begun to reach about the ground state of the doped 2D Hubbard model, with  (?) and without the next nearest neighboring hopping  (?).
@@
Figure 1: Application, architecture, optimization, and performance of Neural Quantum States (NQS) on the Hubbard model. (A) Guided by the variational principle, NQS are powerful and versatile, applicable to a wide range of quantum systems, including ab initio quantum chemistry, spin models, and correlated fermionic models. (B) Various neural network architectures can be used to construct NQS, with increasing expressivity from Restricted Boltzmann Machines (RBM)  (?), Multi-Layer Perceptrons (MLP)  (?, ?, ?, ?) to Transformers. These networks are optimized using methods inspired by quantum information, including Min-SR  (?, ?), SPRING  (?), and the advanced MARCH method proposed in this work. (C) A performance comparison of different NQS methods on the 2D Hubbard model, plotting simulation accuracy against system size. We benchmark the accuracy on systems where exact solutions are available (see Supplementary Section 4.5 for details). This work achieves higher accuracy on larger system sizes than previously reported in other notable NQS studies, demonstrating a significant advancement in the field.
@@
On another front, researchers are also striving to develop more accurate and efficient new methods, aiming to provide more reliable results for strongly correlated systems like the Hubbard model to unveil the underlying exotic physics. In recent years, benefiting from the rapid development in the fields of neural networks and machine learning, Neural Quantum States (NQS) have emerged as a promising new approach for solving quantum many-body systems. Significant progress has been made in the development of advanced NQS methods in quantum systems since their inception by Carleo et al.  (?), which have markedly improved the accuracy of NQS approaches, establishing them as a powerful framework for investigating systems in quantum chemistry  (?, ?, ?) and interacting spin models  (?, ?, ?), as shown in Figure 1(A). In light of these significant achievements, a crucial question arises: how effective is this new paradigm for the study of the more challenging strongly correlated fermionic lattice models?
@@
In this direction, extensive effort has been made to investigate the utilization of NQS methods in the context of the Hubbard model, particularly in the most challenging 2D lightly doped cases. Early attempts to apply NQS approaches to the 2D Hubbard model (?) faced challenges in accurately representing fermionic nodal structures. Subsequent work by Luo et al. (?) introduced a neural network backflow ansatz that substantially improved the representation power of the wavefunction, demonstrating its effectiveness on systems as large as $12\times 8$. More recently in 2022, Robledo et al.  (?) achieved highly accurate results for various geometries of the Hubbard model up to $8\times 8$ by combining Hidden-Fermion Determinant States (HFDS) with Stochastic Reconfiguration (SR)  (?, ?, ?). Despite this notable progress, scaling NQS simulations to even larger, more physically relevant parameter regimes is hindered by network architectural limitations and optimization barriers (?, ?).
@@
In this work, we make significant progress in expanding the scalability and accuracy of the NQS approach for solving the Hubbard model with two key innovations as shown in Figure 1(B). First, a novel NQS architecture is proposed, which leverages the self-attention mechanism  (?), a technique proven to be very effective in scaling up  (?). The built-in long-range structure makes the NQS ansatz different from tensor network states, which consist of only local tensors. This ensures that the NQS can precisely capture long-range correlations and entanglements in strongly correlated systems and can accurately handle larger systems, which is infeasible for other neural network architectures like Restricted Boltzmann Machine (RBM)  (?) and Multi-Layer Perceptron (MLP)  (?, ?, ?, ?). Second, we propose an enhanced optimization algorithm, termed Moment-Adaptive ReConfiguration Heuristic (MARCH), designed to achieve substantially faster and more stable convergence than previous optimization algorithms such as SR  (?, ?, ?), Min-SR  (?, ?) and SPRING  (?), especially for large and frustrated systems.
@@
With these architectural and algorithmic advancements, we achieve state-of-the-art variational ground state energy of the 2D Hubbard model at unprecedented sizes up to $16\times 16$, for both open and periodic boundary conditions, as shown in Figure 1(C). As a result, we confirm the filled stripe ground state in the pure Hubbard model and identify a half-filled stripe ground state with the next nearest neighboring hopping $t^{\prime}=-0.2$, consistent with experimental observations in cuprate  (?). Moreover, in the $t^{\prime}=-0.2$ case, we find the stripe is horizontal, i.e., stripe arrangements prefer the longer side of the geometry. This is quite different from previous calculations on cylinders  (?), indicating the effectiveness of our method in eliminating the boundary effects. We also have a careful analysis of the structure of our NQS ansatz and find that different attention heads can directly encode correlations at different scales after optimization, making it capable to capture long-range correlations and entanglements in the Hubbard model. Our developments in architectures and optimization strategies establish NQS as a powerful and transformative paradigm for solving the Hubbard model. This framework extends readily to a broader range of strongly correlated systems, offering a new lens to explore new exotic quantum many-body states resulting from strong correlations.
@@
Here $\hat{c}_{i\sigma}^{\dagger}$ $(\hat{c}_{j\sigma})$ is the creation (annihilation) operator for an electron with spin $\sigma$ at site $i(j)$, $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$ is the particle number operator. The terms $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote pairs of nearest neighboring and next nearest neighboring sites, respectively. We study the 2D Hubbard model defined on square lattices with size $L_{x}\times L_{y}$. Both open boundary conditions (OBC) and periodic boundary conditions (PBC) are adopted in this work. We set the nearest neighboring hopping amplitude $t=1$ as the unit of energy. The on-site Coulomb repulsion is fixed at $U=8$, a value relevant to cuprates  (?, ?). We consider both the pure Hubbard case ($t^{\prime}=0$) and the case with the next nearest neighboring hopping ($t^{\prime}=-0.2$). In this work, we focus on the lightly doped regime with hole concentration of $\delta=1/8$, which is the most intriguing region in both the 2D Hubbard model and in cuprates.
@@
Figure 2 presents an overview of our NQS algorithm. As shown in Figure 2(A), we employ a transformer-based NQS, parameterized by $\theta$, to represent the amplitude $\psi_{\theta}(\mathbf{n})$ associated with each configuration |$\ket{\mathbf{n}}$. Our transformer-based NQS assigns a unique high-dimensional embedding vector to each site $i$. These embeddings are subsequently processed through a series of multi-head attention blocks, the core component of modern large language models  (?). The attention block is defined as:
@@
where $Q,K,V\in\mathbb{R}^{N\times d_{H}}$ are feature matrices produced by neural networks. Here, $N$ represents the number of sites in the Hubbard model, $d_{H}$ is the hidden dimension of the attention block. The $i$-th row of these matrices corresponds to the $i$-th site in the Hubbard model, arranged in dictionary order. The normalized attention scores $S_{\text{attn}}$, given by $\text{softmax}(QK^{\top}/\sqrt{d_{H}})$, allow the block to aggregate information from all sites by assigning weights. Thus, the attention block can capture complex, long-range correlations, which is crucial for modeling strongly correlated systems like the Hubbard model. The multi-head attention block employs distinct feature matrices to capture diverse correlation patterns in the system. After self-attention layers, a linear layer is employed to produce backflow orbitals $M(\mathbf{n})$. The amplitude is then derived as the determinant of orbitals occupied by electrons.
In our work, we aim to find the ground state by minimizing the total energy of the wavefunction. In previous NQS studies  (?), SR-based optimizers have been used to minimize the energy. These optimizers effectively simulate the imaginary time evolution, enhancing the convergence rate over the standard first-order optimizer by exploiting the quantum geometric tensor  (?). One particular method is the SPRING algorithm  (?), which leverages the history of gradients to accelerate convergence and smooth out the optimization path. Building on this, we introduce the MARCH algorithm. MARCH enhances SPRING by also incorporating an estimate of the second moment of the gradients, drawing a direct parallel to how the popular Adam optimizer  (?) improves upon Momentum  (?). This addition of second-moment information is crucial for navigating the complex and high-dimensional energy landscapes that typically exist in quantum systems. It provides an estimate of the gradient’s variance, allowing MARCH to adapt the learning rate for each parameter individually. This adaptive capability offers a significant advantage, particularly when encountering challenging regions like saddle points or flat plateaus where the gradient is small. While a momentum-based method like SPRING might overshoot, MARCH can use the second-moment information to maintain progress and “march” confidently out of these areas. As illustrated in Figure 2(B), this results in a more direct and efficient path to the energy minimum, leading to faster and more robust convergence.
@@
We then benchmark the new NQS method on the doped pure Hubbard model under OBC with varying sizes $L_{x}$ and $L_{y}$. Direct comparisons of ground-state energies with state-of-the-art DMRG and PEPS  (?) results are presented in Figure 2(C) (with Hartree-Fock energies as reference). In narrow systems such as $L_{y}\leq 6$, the DMRG energy is the lowest. However, when reaching the 2D limit by increasing the width of the system, where DMRG fails to capture the area law of entanglement entropy, the NQS results surpass DMRG. Noticeably, in the entire range of $L_{y}$ we studied, NQS achieves lower variational energies than those reported by state-of-the-art PEPS with bond dimension $D\geq 20$  (?).
@@
Figure 3: Visualization of the attention mechanism. Shown is how a lattice site attends to another target site in the 2D pure Hubbard model for lattice sizes of $8\times 8$, $10\times 10$, and $12\times 12$ at half filling. The four attention heads, which are equivalent and have been rearranged by pattern for clarity, each correspond to a specific physical aspect: short-range correlations, nearest-neighbor hopping, antiferromagnetic correlations, and complex and delocalized correlations in the ground state.
@@
To gain deeper insight into how the transformer architecture captures the intricate physics of the Hubbard model, we examine the attention patterns learned by the network  (?). In our transformer ansatz, the heads after optimization are found to specialize in capturing the correlations in different scales present in the ground-state wave function. This emergent specialization is analogous to observations in fields like natural language processing, where different attention heads in a transformer learn distinct, interpretable patterns such as syntactical dependencies  (?). In Figure 3, we visualize the normalized attention scores $S_{\text{attn}}$ for the half-filling systems, which quantify the learned pairwise importance between sites, from the first layer for representative systems. Similar results for the doped cases or deeper layers can be found in Figures S1-S10 in the Supplementary Materials. The results are striking: the first head clearly learns to prioritize local interactions, with $S_{\text{attn}}$ that decay with the distance between sites, capturing short-range correlations. The second head specializes in nearest-neighbor hopping, with its attention focusing almost exclusively on the four adjacent sites, which is a fundamental process in the Hubbard model. The third head successfully identifies the emergent antiferromagnetic order, assigning high attention scores in a checkerboard pattern characteristic of this magnetic correlation. The final head appears to capture more complex, delocalized phase information, exhibiting a sophisticated pattern that is crucial for accurately representing the full complexity of the wave function. This specialization demonstrates the remarkable ability of the transformer’s attention mechanism to autonomously discover and disentangle the fundamental physical correlations governing the system.
@@
Figure 4: Hole and spin density distributions in the ground state for different lattice sizes with $U=8$ and hole doping $\delta=1/8$ under PBC in the Hubbard model. The magnitudes of the spin density are represented by the sizes of the arrows while the direction is denoted by the direction of the arrows. Hole density is depicted using a color scale. Stripes can be clearly seen in all three systems. (A) Results for $16\times 16$ pure Hubbard model where the stripe state with $\lambda=8$ is observed. (B) Results for $16\times 12$ system with $t^{\prime}=-0.2$ where the stripe state with $\lambda=4$ is observed. (C) Results for $32\times 8$ system with $t^{\prime}=-0.2$ where the stripe state with $\lambda=4$ is observed. Notice that the stripes in the ground state for $t^{\prime}=-0.2$ in (B) and (C) are horizontal, even though both vertical and horizontal stripes are allowed in the systems. More results can be found in the supplementary materials.
The ground state of the lightly doped pure ($t^{\prime}=0$) 2D Hubbard model was established as the filled stripe before  (?). Recently Liu et al.  (?) observed a crossover of the stripe direction with the increase of system widths for systems under OBC, indicating the existence of a large finite size effect in systems with open boundaries. We performed systematic calculations of the pure Hubbard model for different sizes under both OBC and PBC. For systems with OBC, we find stripe states that match well with the results obtained from PEPS but with lower energies (see Figure 2(C)). In particular, our results also demonstrate a crossover between horizontal and vertical stripes as a function of system width (from $L_{y}=8$ to $L_{y}=12$, see Figures S14-S15 in the Supplementary Materials). But when we switch to PBC, the crossover vanishes (see Figures S17-S18 in the Supplementary Materials), indicating the crossover is a finite size and open boundary effect. The ability to handle systems under PBC with NQS can mitigate the finite size effect, allowing to obtain more reliable results in the thermodynamic limit. By studying systems under PBC, we observe a robust, filled stripe phase with wavelength of $\lambda=8$ as shown in Figure 4(A), a result that is consistent with previous findings  (?). Simulations on smaller $8\times 8$ and $12\times 12$ lattices (see Figures S19-S20 in the Supplementary Materials) show clear finite-size effects due to their incommensurability with the wavelength $\lambda=8$.
@@
It was found that the doped pure Hubbard model doesn’t host superconductivity  (?). To pursue superconducting ground states, the simplest modification is to include the next nearest neighboring hopping ($t^{\prime}$) term. On one hand, a finite $t^{\prime}$ can account for the electron and hole doping asymmetry in the phase diagram of cuprates. On the other hand, electronic structure calculations  (?, ?) show a finite $t^{\prime}\approx-0.2$ term indeed exists in the one band effective Hamiltonian of cuprates. A recent study by Xu et al.  (?) found that the phase diagram of the Hubbard model with $t^{\prime}=-0.2$ agrees well with the cuprates, where fractional filling stripe coexists with the d-wave superconductivity in the hole doping region. The $t^{\prime}$ term frustrates the anti-ferromagnetic order in the parent compound and makes the doped stripe state fragile. Importantly, large finite size and boundary effect  (?) was found for the wavelength of the stripe state in the hole doping region, which fluctuates in the vicinity of $\lambda=4$  (?).
We perform systematic calculations for systems under PBC with $t^{\prime}=-0.2$. With PBC, we can get rid of the boundary effect. Our calculations on large sizes show the ground state for the $t^{\prime}=-0.2$ Hubbard model is clearly the half-filled stripe, i.e., with wavelength $\lambda=4$, as shown in Figure 4(B). These results indicate that the previous found fluctuating wavelengths for stripe may result from the finite size and boundary conditions. The $\lambda=4$ stripe agrees well with the experimental observation of cuprates  (?), providing further evidence to support the relevance of the $t^{\prime}$ Hubbard model to cuprates.